PIQ (Pritchard’s Intelligent Quotient) scores are normally distributed with a mean of 120 and a standard deviation of 10.4.

a. What is the probability that a person selected at random has a PIQ score greater than 135?

b. What is the probability that an individual has a PIQ score between 104 and 126?

c. The gifted program at Centereach High School accepts students with a PIQ score in the top 10% of the population. What PIQ score would a student need to have to be accepted to the program?

d. Find the PIQ score at the 55th percentile.
Answer #1

a)

for normal distribution z score =(X-\mu)/\sigma
here mean=      \mu= 120
std deviation   =\sigma= 10.400

  probability that a person selected at random has a PIQ score greater than 135:

probability = P(X>135) = P(Z>1.44)= 1-P(Z<1.44)= 1-0.9251= 0.0749

b)

  probability that an individual has a PIQ score between 104 and 126:

probability = P(104<X<126) = P(-1.54<Z<0.58)= 0.7190-0.0618= 0.6572

c)

for highest 10 percentile ; critical z =1.28

hence corresponding score =mean +z*std deviation =120+1.28*10.4= 133.3~ 134

d)

for 55th percentile ; critical z =0.13

hence corresponding score =mean +z*std deviation =120+0.13*10.4= 121.35